# thanks: ORCID $0$-0001-9980-0363thanks: ORCID $0$-0001-9896-1018thanks: ORCID $0$-0003-1168-7883

### v1

@@ IV Discussion
where $E=2\sigma_{39}\sigma_{40}+2\sigma_{50}\sigma_{1}$. Up to the last term $E$, $Z_{-}$ is also equal to the partition function of the one-dimensional Ising model. This term introduces some finite size corrections which can be neglected when the size $n\times n$ of the lattice increases $n\to\infty$, so we see that for large lattices the original model is mapped into a one-dimensional Ising model on a chain of length $2n$. Energy per node for the Ising chain is known to be $-\tanh\beta$, so therefore in the Heider model with asynchronous update scheme energy per triangle should reproduce this result $U/\Delta\rightarrow-\tanh\beta$ in the limit $n\to\infty$. This is not the case for synchronous update scheme which violates the detailed balance condition.
@@ Acknowledgements.
The authors thank Zdzisław Burda for a fruitful discussion of the gauge invariance theory applied to show that the perfectly balanced state is unreachable. We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016259.
@@ Appendix A Archimedean lattices mapped into square lattice and the shapes of their adjacency matrices
In Figures 13(a), 10(a), 11(a), 12(a) and 9(a) shapes of Archimedean lattices embedded in a distorted triangular lattice are presented for kagomé, $(3^{2},4,3,4)$, $(3^{3},4^{2})$, maple-leaf, and triangular lattices, respectively.
@@
In Figures 13(b), 10(b), 11(b), 12(b) and 9(b) examples of adjacency matrices associated with lattices presented in Figures 13(a), 10(a), 11(a), 12(a) and 9(a) are displayed.
@@
Figure 9: Triangular $(3^{6})$ lattice, $k=6$, (9(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=6$ (9(b)) and its adjacency matrix
@@
Figure 10: Archimedean $(3^{2},4,3,4)$ lattice, $k=5$, (10(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=4$ (10(b)) and its adjacency matrix
@@
Figure 11: Archimedean $(3^{3},4^{2})$ lattice, $k=5$, (11(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=6$ (11(b)) and its adjacency matrix
@@
Figure 12: Maple leaf $(3^{4},6)$ lattice, $k=5$, (12(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=7$ (12(b)) and its adjacency matrix
@@
Figure 13: Archimedean $(3,6,3,6)$ lattice (kagomé lattice), $k=4$, (13(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=6$ (13(b)) and its adjacency matrix
Appendix B Listings
LABEL:lst:module-par presents Fortran module defining constant parameters for the subroutines supplying the adjacency matrix $\mathbf{A}$. The integer variable Z corresponds to the network connectivity $k$. The integer variable W represents the linear size of the underlying square lattice in which the distorted triangles are embedded. Then, the integer variable N is equal to the square of W. To ensure the possibility of wrapping the lattice according to the assumed periodic boundary condition, its value must be a multiple of 4 for $(3^{2},4,3,4)$ and 7 for $(3^{4},6)$. The other lattices are presented for W=6, however for $(3^{3},4^{2})$ and $(3,6,3,6)$ any even number is suitable, while there are no restrictions in the case of the triangular lattice $(3^{6})$.
LABEL:lst:code-triangular, LABEL:lst:code-al33344, LABEL:lst:code-tri2kagome, LABEL:lst:code-tri2mapleleaf and LABEL:lst:code-tri2al33434 contain Fortran subroutines computing the adjacency matrix $\mathbf{A}$ for triangular, $(3^{2},4,3,4)$, $(3^{3},4^{2})$, $(3^{4},6)$, and $(3,6,3,6)$, respectively. Subroutines tri_2_al33434 (LABEL:lst:code-tri2al33434), tri_2_maple_leaf (LABEL:lst:code-tri2mapleleaf), tri_2_kagome (LABEL:lst:code-tri2kagome) for the $(3^{2},4,3,4)$, $(3^{3},6)$, $(3,6,3,6)$ lattices, respectively, require earlier preparation of adjacency matrix $\mathbf{A}$ as for triangular lattices using the subroutine given in LABEL:lst:code-triangular. The adjacency matrix $\mathbf{A}$ is returned in the matrix variable A. The local matrix variable nn (if used) stores labels of $k$ neighbors of the sites in the lattice.
Listing 1: Module including parameters
1MODULE par
2INTEGER, PARAMETER :: Z=5, W=7, N=W*W
3END MODULE
Listing 2: Subroutine for triangular $(3^{6})$ lattice
1SUBROUTINE triangular(A)
2!! -------------------------------------------
3USE par
4IMPLICIT NONE
5INTEGER, DIMENSION(0:N-1,0:Z-1) :: nn
6INTEGER, DIMENSION(0:N-1,0:N-1) :: A
7INTEGER :: i, j
8!! -------------------------------------------
9DO i=0, N-1
10 nn(i,0) = MOD(N+i -W ,N)
11 nn(i,1) = MOD(N+i -W+1,N)
12 nn(i,2) = MOD(N+i -1 ,N)
13 nn(i,3) = MOD(N+i +1 ,N)
14 nn(i,4) = MOD(N+i +W-1,N)
15 nn(i,5) = MOD(N+i +W ,N)
16 IF(MOD(i,W)==0) THEN
17 nn(i,2) = MOD(N+W+i -1 ,N)
18 nn(i,4) = MOD(N+W+i +W-1,N)
19 ENDIF
20 IF(MOD(i+1,W)==0) THEN
21 nn(i,1) = MOD(N-W+i -W+1,N)
22 nn(i,3) = MOD(N-W+i +1 ,N)
23 ENDIF
24ENDDO
25!! -------------------------------------------
26A=0
27DO i=0,N-1
28DO j=0,Z-1
29 A(i,nn(i,j))=1
30ENDDO
31ENDDO
32!! -------------------------------------------
33END SUBROUTINE triangular
Listing 3: Subroutine for Archimedean $(3^{2},4,3,4)$ lattice
1SUBROUTINE tri_2_al33434(A)
2USE par
3IMPLICIT NONE
4INTEGER, DIMENSION(0:N-1,0:Z-1) :: nn
5INTEGER, DIMENSION(0:N-1,0:N-1) :: A
6INTEGER :: i, x ,y
7!! -------------------------------------------
8!! neighbors on triangular lattice
9DO i=0, N-1
10 nn(i,0) = MOD(N+i -W ,N)
11 nn(i,1) = MOD(N+i -W+1,N)
12 nn(i,2) = MOD(N+i -1 ,N)
13 nn(i,3) = MOD(N+i +1 ,N)
14 nn(i,4) = MOD(N+i +W-1,N)
15 nn(i,5) = MOD(N+i +W ,N)
16 IF(MOD(i,W)==0) THEN
17 nn(i,2) = MOD(N+W+i -1 ,N)
18 nn(i,4) = MOD(N+W+i +W-1,N)
19 ENDIF
20 IF(MOD(i+1,W)==0) THEN
21 nn(i,1) = MOD(N-W+i -W+1,N)
22 nn(i,3) = MOD(N-W+i +1 ,N)
23 ENDIF
24ENDDO
25!! -------------------------------------------
26!! adjacency matrix for (3^2,4,3,4) lattice:
27DO x=0,W-1
28DO y=0,W-1
29 i=x+y*W
30 IF(MOD(x,W).EQ.0 .AND. MOD(y,W).EQ.0)
31 A(i,nn(i,1))=0
32 IF(MOD(x,W).EQ.1 .AND. MOD(y,W).EQ.0)
33 A(i,nn(i,5))=0
34 IF(MOD(x,W).EQ.2 .AND. MOD(y,W).EQ.0)
35 A(i,nn(i,1))=0
36 IF(MOD(x,W).EQ.3 .AND. MOD(y,W).EQ.0)
37 A(i,nn(i,5))=0
38
39 IF(MOD(x,W).EQ.0 .AND. MOD(y,W).EQ.1)
40 A(i,nn(i,4))=0
41 IF(MOD(x,W).EQ.1 .AND. MOD(y,W).EQ.1)
42 A(i,nn(i,0))=0
43 IF(MOD(x,W).EQ.2 .AND. MOD(y,W).EQ.1)
44 A(i,nn(i,4))=0
45 IF(MOD(x,W).EQ.3 .AND. MOD(y,W).EQ.1)
46 A(i,nn(i,0))=0
47
48 IF(MOD(x,W).EQ.0 .AND. MOD(y,W).EQ.2)
49 A(i,nn(i,5))=0
50 IF(MOD(x,W).EQ.1 .AND. MOD(y,W).EQ.2)
51 A(i,nn(i,1))=0
52 IF(MOD(x,W).EQ.2 .AND. MOD(y,W).EQ.2)
53 A(i,nn(i,5))=0
54 IF(MOD(x,W).EQ.3 .AND. MOD(y,W).EQ.2)
55 A(i,nn(i,1))=0
56
57 IF(MOD(x,W).EQ.0 .AND. MOD(y,W).EQ.3)
58 A(i,nn(i,0))=0
59 IF(MOD(x,W).EQ.1 .AND. MOD(y,W).EQ.3)
60 A(i,nn(i,4))=0
61 IF(MOD(x,W).EQ.2 .AND. MOD(y,W).EQ.3)
62 A(i,nn(i,0))=0
63 IF(MOD(x,W).EQ.3 .AND. MOD(y,W).EQ.3)
64 A(i,nn(i,4))=0
65ENDDO
66ENDDO
67END SUBROUTINE tri_2_al33434
Listing 4: Subroutine for Archimedean $(3^{3},4^{2})$ lattice
1SUBROUTINE al_33344(A)
2!! -------------------------------------------
3USE par
4IMPLICIT NONE
5INTEGER, DIMENSION(0:N-1,0:Z-1) :: nn
6INTEGER, DIMENSION(0:N-1,0:N-1) :: A
7INTEGER :: i, j, x, y
8!! -------------------------------------------
9DO y=0,W-1,2
10DO x=0,W-1
11 i=x+y*W
12 nn(i,0) = MOD(N+i -1 ,N)
13 nn(i,1) = MOD(N+i +1 ,N)
14 nn(i,2) = MOD(N+i -W ,N)
15 nn(i,3) = MOD(N+i +W ,N)
16 nn(i,4) = MOD(N+i +W-1,N)
17 IF(MOD(i,W)==0) THEN
18 nn(i,0) = MOD(N+W+i -1 ,N)
19 nn(i,4) = MOD(N+W+i +W-1,N)
20 ENDIF
21 IF(MOD(i+1,W)==0) THEN
22 nn(i,1) = MOD(N-W+i +1,N)
23 ENDIF
24ENDDO
25ENDDO
26!! -------------------------------------------
27DO y=1,W-1,2
28DO x=0,W-1
29 i=x+y*W
30 nn(i,0) = MOD(N+i -1 ,N)
31 nn(i,1) = MOD(N+i +1 ,N)
32 nn(i,2) = MOD(N+i -W ,N)
33 nn(i,3) = MOD(N+i +W ,N)
34 nn(i,4) = MOD(N+i -W+1,N)
35 IF(MOD(i,W)==0) THEN
36 nn(i,0) = MOD(N+W+i -1,N)
37 ENDIF
38 IF(MOD(i+1,W)==0) THEN
39 nn(i,1) = MOD(N-W+i +1,N)
40 nn(i,4) = MOD(N-W+i -W+1,N)
41 ENDIF
42ENDDO
43ENDDO
44!! -------------------------------------------
45A=0
46DO i=0,N-1
47DO j=0,Z-1
48 A(i,nn(i,j))=1
49ENDDO
50ENDDO
51!! -------------------------------------------
52END SUBROUTINE al_33344
Listing 5: Subroutine for maple leaf $(3^{3},6)$ lattice
1SUBROUTINE tri_2_maple_leaf(A)
2!! -------------------------------------------
3USE par
4IMPLICIT NONE
5INTEGER, DIMENSION(0:N-1,0:N-1) :: A
6INTEGER :: i, j, k, x, y
7!! -------------------------------------------
8DO y=0,W-1
9DO k=1,S
10 x = MOD(2*y+7*k,W)
11 i = x + W*y
12 DO j=0,N-1
13 A(i,j)=0
14 A(j,i)=0
15 ENDDO
16ENDDO
17ENDDO
18!! -------------------------------------------
19END SUBROUTINE tri_2_maple_leaf
Listing 6: Subroutine for kagomé $(3,6,3,6)$ lattice
1SUBROUTINE tri_2_kagome(A)
2!! -------------------------------------------
3USE par
4IMPLICIT NONE
5INTEGER, DIMENSION(0:N-1,0:N-1) :: A
6INTEGER :: i, j, x, y
7!! -------------------------------------------
8DO y=1,W,2
9DO x=1,W,2
10 i=x+W*y
11 DO j=0,N-1
12 A(i,j)=0
13 A(j,i)=0
14 ENDDO
15ENDDO
16ENDDO
17!! -------------------------------------------
18END SUBROUTINE tri_2_kagome

### v2

@@ III.1 Calculations
where $E=2\sigma_{39}\sigma_{40}+2\sigma_{50}\sigma_{1}$. Up to the last term $E$, $Z_{-}$ is also equal to the partition function of the one-dimensional Ising model. This term introduces some finite size corrections which can be neglected when the size $N=n\times n$ of the lattice increases with $n\to\infty$, so we see that for large lattices the original model is mapped into a one-dimensional Ising model on a chain of length $2n$. The mean energy per node for the Ising chain is known to be $-\tanh\beta$, so therefore in the Heider model with the asynchronous update scheme, the mean energy per triangle should reproduce this result $U/\Delta\to-\tanh(\beta)$ in the limit $N\to\infty$.
III.2 Computations
[…]
Figure 7: $\langle U_{\text{av}}\rangle/\Delta$ for synchronous (7(a), 7(c), 7(e), 7(g)), and asynchronous (7(b), 7(d), 7(f), 7(h)) update scheme. Nodes degree (on average, for the diluted triangular lattice and classical random graphs) $k=4$ (7(c), 7(d)), $k=5$ (7(e), 7(f)), $k=6$ (7(g), 7(h))
The mean energy per triangle defined in Equation 3 is therefore calculated in the final 90% of the time steps and then averaged to find $U_{\text{av}}$, and then averaged again over all the simulations performed. The results, denoted as $\langle U_{\text{av}}\rangle/\Delta$ are presented in LABEL:{fig:U_vs_T} as a function of temperature $T$, together with the curve given by
$U/\Delta=-\tanh(1/T),$ (24)
corresponding to the theoretical results of Section III.1. For all simulations on $(3^{2},4,3,4)$ and $(3^{3},4^{2})$ lattices we assume $N=64$ which corresponds to $\Delta=64$; on $(3^{4},5)$ $N=49$, $\Delta=56$; on the triangular lattice $(3^{6})$ $N=49$, $\Delta=98$; and on $(3,6,3,6)$ $N=100$ with $\Delta=50$.
For comparison, we match the Archimedean lattices with the corresponding diluted triangular lattices (DTL) [14] and classical random graphs (CRG) [16] with $N=64$ and having the same average node degree $k=4$, $5$, or $6$ as the connectivity of the considered Archimedean lattices. The only Archimedean lattice discussed here with $k=4$ is $(3,6,3,6)$ and for synchronous updating DTL and CRG of the same $k$ reveal similar but not identical characteristics, with non-perfect balance $\langle U_{\text{av}}(T)\rangle/\Delta\approx-0.9$ at $T\to 0$ as visible in Figure 7(c). It should be noted, though, that the random nature of the DTL and CRG graphs means that the simulations are performed on lattices with the same average node degree and number of nodes, but with different structure of connections (graph egdes) which results, among other things, in the number of triads in those lattices that is not constant in all simulations but found to be $\Delta=37\pm 3$ for DTL or $\Delta=11\pm 3$ for CRG, with $N=64$ in both cases when $k=4$. For $k=5$ (with $\Delta=75\pm 3$ for DTL or $\Delta=20\pm 4$ for CRG, Figure 7(e)) and $k=6$ (with $\Delta=36\pm 6$ for CRG, Figure 7(g)) with synchronous updating the results for DTL and CRG are similar to those for $k=4$, which also means that they are not exactly in line with those for Archimedean lattices, in particular in the case of $(3^{2},4,3,4)$ lattice where the balanced state is not observed even in the low-temperature limit. At higher $T$, the results for all Archimedean, DTL and CRG lattices agree. For asynchronous updating, the results obtained for DTL and CRG at $k=4,5,6$ and presented in Figures 7(d), 7(f) and 7(h) do not differ from those of Archimedean lattices.
Figure 8: Values of $U_{\text{av}}/\Delta$ obtained in 100 simulations for each temperature $T$, for various simulation times $t_{\text{max}}=10^{4}$, $10^{6}$ and $10^{8}$, for (8(a)) $(3^{4},6)$ and (8(b)) $(3^{2},4,3,4)$ lattice and synchronous scheme of link updates
The dispersion of the values obtained in 100 simulations is usually relatively small, especially when asynchronous updating is applied. In most cases, it is also true for synchronous updating, except for the $(3^{2},4,3,4)$ and $(3^{4},6)$ lattices at $T<1$ where it was required to perform $t_{\text{max}}=10^{8}$ time steps, and for DTL and CRG lattices where $t_{\text{max}}=10^{6}$ was used. In particular in the case of $(3^{4},6)$ lattice shorter simulations do not allow us to see that the balanced state is achieved at $T\to 0$, as visible in Figure 8(a) showing the values $U_{\text{av}}/\Delta$ obtained in all performed simulations. Figure 8(b) confirms that on the $(3^{2},4,3,4)$ lattice the balanced state is not observed even in the low-temperature limit, and additionally the values of the work function obtained in the simulations performed are quantized and distributed over a relatively long interval. The values of the work function observed at the lowest $T$ are spaced in the intervals of $2/\Delta$, which corresponds to the difference related to the energetically nearest states. The same quantization of the work function applies for any other planar lattice that contains triangles as it is associated with changing a single term ($x_{ij}x_{jk}x_{ik}$—where $i$, $j$, $k$ are triangle vertices) in sum in nominator of Equation 3 from $+1$ to $-1$ (or vice versa).
Figure 9: Thermal evolution (in the low temperature limit $T<1$) of fractions of triads presented in Figure 1 in the case of synchronous updating and different Archimedean lattices. The panels (9(a)), (9(b)), (9(c)), and (9(d)) correspond to the triangles presented in Figures 1(a), 1(b), 1(c), and 1(d), respectively. The error bars show the standard deviation of the results in 100 simulations. Dashed line for asynchronous updating (common result for all lattices)
If any of the systems discussed is in the balanced state, it means that all triads are balanced, that is, they are in one of two possible configurations: with all positive links ([$+$][$+$][ + + + ], Figure 1(a)) or with one positive and two negative links ([+][+--][ + - - ], Figure 1(c)). The fraction of triads of a given kind, $\langle f\rangle$, as a function of $T$ is shown in Figure 9. It reveals that in the cases where the balanced state is reached at $T\to 0$, about 25% of the triads are in the [$+$][$+$][ + + + ] configuration, and the rest are in the [+][+--][ + - - ] state. However, in that limit (only for $T<0.2$) those fractions vary substantially among the simulations, as shown by the error bars indicating the standard deviation, which is also the case for the $(3^{2},4,3,4)$ lattice which does not reach the balanced state and for all triads in that case including [++][++-][ + + - ] (Figure 1(b)) and [][---][ - - - ] (Figure 1(d)). On the other hand, the $(3^{6})$ triangular lattice produces almost identical results in all simulations. In the balanced state observed in the low-temperature limit, the fractions of triads obtained using the asynchronous updating scheme (dashed lines in Figure 9) are the same as for the balanced state reached with synchronous updating. Yet, in the case of asynchronous updating the increasing temperature causes much slower decrease of the number of balanced triads than for the synchronous approach, which is also visible in slower increase of the work function with increasing $T$.
IV Discussion
We could expect that the results obtained with synchronous updating give systematically larger values of $U(T)$, than the thermodynamical counterparts. Further, and for the same reason, the frequencies of imbalanced triads obtained for the synchronous updating should exceed those obtained within the asynchronous scheme. The results from the synchronous updating are expected to depend on the length of simulation time because the system can escape from local minima of the work function after some characteristic time, which should depend on temperature. Our numerical results, shown in Figure 8, indicate that this effect appears indeed. Moreover, it is different in different Archimedean lattices. Apparently, the structure of local minima of the work function is sensitive to the lattice structure.
Accordingly, the synchronous updating for lower $T$ produces results that differ significantly from one Archimedean lattice to another. Figure 7(a) shows that the synchronous updating leads to $\langle U_{\text{av}}(T)\rangle/\Delta$ dependencies which are unique to each of the discussed Archimedean lattices. In the low-temperature limit, we observe perfectly balanced states, $\langle U_{\text{av}}\rangle/\Delta=-1$, in the $(3,6,3,6)$, $(3^{3},4^{2})$ and $(3^{4},6)$ lattices, but the decrease of energy with decreasing temperature occurs along a different curve for each of those lattices, and in all cases the values of the work function per triangle remain significantly larger than $-\tanh(1/T)$. On the other hand, only a slight decrease in the work function is found for the triangular lattice $(3^{6})$ and somewhat larger, to $\langle U_{\text{av}}(T\to 0)\rangle/\Delta\approx-0.5$, for the $(3^{2},4,3,4)$ lattice. For the asynchronous updating, the results in Figure 7(b) prove that for all types of the considered Archimedean lattices the temperature dependence of the work function per triad is exactly the same and identical to the theoretically predicted $U/\Delta=-\tanh(1/T)$.
Besides local minima of $U$ and the related metastable (‘jammed’ [35]) states, the synchronous updating at $T=0$ admits limit cycles. In particular, limit cycles of length two, three and more have been identified in Reference 40. In the case of limit cycles of at least three steps, the detailed balance condition [41, p. 406] is violated. To see this, consider three states $A$, $B$ and $C$ in a limit cycle: $A$ is switched to $B$, $B$ to $C$ and $C$ to $A$. Their stationary probabilities are equal, but $B$ is never switched to $A$. For $T>0$, the transitions are not deterministic, yet the asymmetry of the rates of switching remains. We admit, however, that for positive temperature, the contribution of limit cycles to the system behavior is hard to be evaluated.
The equivalence of the planar Heider system and the one-dimensional Ising model, demonstrated at the end of Section III.1, allows us to deduce the lack of balance-imbalance phase transition at $T>0$. It is worthwhile to add that the consequence of this equivalence, i.e. the inequality $U(T)>-1$, seems to hold also for non-planar structures.
As remarked in the Introduction, the synchronous updating scheme finds a counterpart in parallel decisions made by groups of voters. This scheme can also be used in settings of games played in parallel, as for example the Prisoner’s Dilemma [42]. There, the matrix of connections between players can be established by an appropriate structure of channels of exchange of information. Further, the incompleteness of information, characteristic for evolutionary games [42], can be modeled by the thermal noise.
To summarize, the shape of the mean work function dependence on temperature for all planar networks investigated here, including the triangular lattice, is the same as for a set of non-interacting triads. On the contrary, it appears different for a network of tetrahedrons, i.e. cliques of four nodes. The result for one tetrahedron is the same as for a network of tetrahedra, connected only by corners. Moreover, both above functions $U(T)$ are different from the one of a clique of five nodes. Finally, the latter formula for one clique is appropriate also for a network of such cliques, mutually connected by corners.
Concluding, the function $U(T)$ depends on the local topology of the network. All planar structures of triads, including the triangular lattice, display the same thermal dependence of $U$. Further, an interaction of cliques through shared corners is not relevant for $U(T)$. On the contrary, the size of cliques matters. This raises an interesting question of whether and under which conditions the balance-imbalance transition is a collective effect.
@@ Acknowledgements.
The authors thank Zdzisław Burda for a fruitful discussion of the gauge invariance theory applied to show that the perfectly balanced state is unreachable (Section III.1). We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016259.
@@ Appendix A Archimedean lattices mapped into square lattice and the shapes of their adjacency matrices
In Figures 14(a), 11(a), 12(a), 13(a) and 10(a) shapes of Archimedean lattices embedded in a distorted triangular lattice are presented for kagomé, $(3^{2},4,3,4)$, $(3^{3},4^{2})$, maple-leaf, and triangular lattices, respectively.
@@
In Figures 14(b), 11(b), 12(b), 13(b) and 10(b) examples of adjacency matrices associated with lattices presented in Figures 14(a), 11(a), 12(a), 13(a) and 10(a) are displayed.
@@
Figure 10: Triangular $(3^{6})$ lattice, $k=6$, (10(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=6$ (10(b)) and its adjacency matrix
@@
Figure 11: Archimedean $(3^{2},4,3,4)$ lattice, $k=5$, (11(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=4$ (11(b)) and its adjacency matrix
@@
Figure 12: Archimedean $(3^{3},4^{2})$ lattice, $k=5$, (12(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=6$ (12(b)) and its adjacency matrix
@@
Figure 13: Maple leaf $(3^{4},6)$ lattice, $k=5$, (13(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=7$ (13(b)) and its adjacency matrix
@@
Figure 14: Archimedean $(3,6,3,6)$ lattice (kagomé lattice), $k=4$, (14(a)) $N=W^{2}$ nodes at and inside green square are labeled as $i=0,\cdots,W^{2}-1$, $i=x+Wy$, $x,y\in\{0,1,2,\cdots,W-1\}$, $W=6$ (14(b)) and its adjacency matrix